OpenStudy (anonymous):

Evaluate expression for b=3 and c=7 bc square (not sure how to write the square on computer)

OpenStudy (saifoo.khan):

\[(bc)^2 \]or bc^2

OpenStudy (anonymous):

the 1st expression

OpenStudy (anonymous):

(b^3)^2 (c^2)^7?

OpenStudy (saifoo.khan):

\[(3 \times 7 )^2=(21)^2=441\]

OpenStudy (anonymous):

(3 multiply by 7) = (21 ) (square) = 441

OpenStudy (anonymous):

its (3 multiply by 7) (square)